# Stabilizing Extreme Q-learning by Maclaurin Expansion

### v1

@@ Abstract
In Extreme Q-learning (XQL), Gumbel Regression is performed with an assumed Gumbel distribution for the error distribution. This allows learning of the value function without sampling out-of-distribution actions and has shown excellent performance mainly in Offline RL. However, issues remained, including the exponential term in the loss function causing instability and the potential for an error distribution diverging from the Gumbel distribution. Therefore, we propose Maclaurin Expanded Extreme Q-learning to enhance stability. In this method, applying Maclaurin expansion to the loss function in XQL enhances stability against large errors. It also allows adjusting the error distribution assumption from normal to Gumbel based on the expansion order. Our method significantly stabilizes learning in Online RL tasks from DM Control, where XQL was previously unstable. Additionally, it improves performance in several Offline RL tasks from D4RL, where XQL already showed excellent results.
@@ 1 Introduction
Thus, in this study, we propose a both simple and practical algorithm, Maclaurin Expanded Extreme Q-learning (MXQL), which not only stabilizes the Gumbel loss but also allows for the adjustment of the error distribution assumption from normal to Gumbel. The proposed method uses Expanded Gumbel loss, which is a Maclaurin expansion of the Gumbel loss. By reducing the order of expansion $n$, this loss, as shown to the left of Figure 1, mitigates excessively large or small gradients compared to the Gumbel loss. We consider this loss from the perspective of the error distribution. When the order of expansion is high, it naturally approaches the Gumbel loss, and the assumed error distribution approaches the Gumbel distribution. When the order of expansion is two, the loss function becomes equivalent to L2 loss, assuming a normal distribution for error distribution. In other words, by changing the order of expansion, it is possible to adjust the assumed error distribution between normal and Gumbel distributions. The assumed error distribution is illustrated to the right of Figure 1, showing that as $n$ increases, the distribution transitions from normal to Gumbel. This algorithm can be implemented by rewriting just a few lines of the loss function of XQL, significantly improving stability. In the experiments, similar to Garg et al. (2023), we compared performance using DM Control tasks (Tassa et al., 2018; Tunyasuvunakool et al., 2020) for Online RL and D4RL tasks (Fu et al., 2020) for Offline RL. In Online RL scenarios, where XQL was previously unstable, we observed improved stability. Furthermore, in Offline RL, our method demonstrated superior performance compared to existing methods, including XQL, across several tasks.
@@ 2.2.2 Gumbel Regression
In XQL, this property is applied to Q-learning, directly estimating the soft-value function ($V^{*}(s)=\beta\log\sum_{a}\mu(a\mid s)\text{exp}(Q(s,a)/\beta)$) in entropy maximization RL. The loss function becomes:
@@ 2.2.3 Role of $\beta$
The Q-function is updated as:
$Q^{k+1}(s,a)=r(s,a)+\gamma\mathbb{E}[Q^{k}(s^{\prime},a^{\prime})]-\beta\log\frac{\pi(a\mid s)}{\mu(a\mid s)}$ (5)
Here, $\beta\log\pi(a\mid s)/\mu(a\mid s)$ serves as a conservative term against deviations from the behavior policy, with $\beta$ modulating the level of conservatism. As can be seen from eq. 4, this parameter, $\beta$, also impacts learning stability: small $\beta$ values can increase gradients, risking divergence, while large $\beta$ values may reduce gradient magnitudes, decelerating convergence. Therefore, while $\beta$ modifies conservatism, it does not ensure stability across all values.
@@ 3.1 Instability related to $\beta$
For scalar data represented as $x_{i}$, parameter $h$, and Gumbel noise $\epsilon_{i}$, we perform Gumbel Regression as $x_{i}=h+\epsilon_{i}$ where $\epsilon_{i}\sim-\mathcal{G}(0,\beta_{reg}),x_{i}\sim-\mathcal{G}(0,\beta_{data})$. This implies that when $\beta_{reg}$ and $\beta_{data}$ are equal, the assumptions concerning the error distribution in Gumbel Regression are fully met. We varied the $\beta$ of the data ($\beta_{data}$) and the $\beta$ used in Gumbel Regression ($\beta_{reg}$) respectively, estimated them, and analyzed how differences between $\beta_{data}$ and $\beta_{reg}$ affected the estimation results. The estimated $h$ was assessed based on its deviation from the true value ($\log\sum e^{x_{i}/\beta_{reg}}$). The results are presented in Figure 2.
@@ 3.2 Expanded Gumbel Loss
where $n$ is the order of expansion. As shown on the left side of Figure 1, the Expanded Gumbel loss exhibits a less steep gradient on the right side and a steeper gradient on the left side compared to the traditional Gumbel loss. This modification resolves issues such as the collapse of learning and slow convergence observed in the preliminary experiments. Moreover, by adjusting the expansion order $n$, the degree of stabilization can be altered: smaller values of $n$ lead to greater deviation from the Gumbel loss but enhance stability. It should be noted that $n$ is selected from even numbers to ensure the loss remains positive. In MXQL, similar to XQL, the Expanded Gumbel loss is employed for estimating the value function, and the loss function is defined as follows:
@@ 3.3 Perspective of Error Distribution
We name our method, which substitutes XQL’s Gumbel loss with Expanded Gumbel loss, Maclaurin Expanded Extreme Q-learning (MXQL). Our approach is based on XQL, simply changing its loss function. The implementation details are provided below, demonstrating that it requires only minimal changes from XQL.
loss = 0.
for i in range(2, n+1):
loss += (label - pred)**i / (math.factorial(i) * (beta**i))
return loss
In MXQL, the expansion order $n$ is a hyperparameter. However, it should be noted that the clipping size is also a hyperparameter in XQL, and the number of hyperparameters has not changed compared to XQL.
@@ 4.2 Offline RL
To ensure a fair comparison with XQL, experiments were conducted on the same Gym locomotion, AntMaze, and Franka Kitchen tasks as those reported in Garg et al. (2023). The scores are displayed in Table 1 and are compared with those from Brandfonbrener et al. (2021); Fujimoto & Gu (2021); Kumar et al. (2020); Kostrikov et al. (2022); Garg et al. (2023). The proposed method, MXQL, attained average scores that were either superior to or on par with those of existing methods as reported in Kostrikov et al. (2022); Garg et al. (2023) across various environments. Specifically, significant gains were observed in the hopper-medium and kitchen-complete tasks. For the medium and large AntMaze tasks, the results were inferior compared to those of existing methods, with XQL’s stabilization proving more suitable than MXQL’s. Future research is warranted to explore the differences in appropriate stabilization methods depending on the task. The average scores for each domain upon varying $n$ are presented in Table 3. The small fluctuation in scores concerning $n$ suggests that MXQL is relatively insensitive to changes in $n$ within the range of 4 to 12.
@@
halfcheetah-med 42.6 48.4 48.3 44.0 47.4 47.7 47.8 ±0.1
hopper-med 52.9 59.6 59.3 58.5 66.3 71.1 89.7 ±3.6
walker2d-med 75.3 81.8 83.7 72.5 78.3 81.5 86.1 ±0.2
halfcheetah-med-rep 36.6 38.1 44.6 45.5 44.2 44.8 45.7 ±0.3
hopper-med-rep 18.1 97.5 60.9 95.0 94.7 97.3 101.8 ±0.8
walker2d-med-rep 26.0 49.5 81.8 77.2 73.9 75.9 88.8 ±1.5
halfcheetah-med-exp 55.2 93.4 90.7 91.6 86.7 89.8 93.3 ±0.7
hopper-med-exp 52.5 103.3 98.0 105.4 91.5 107.1 113.0 ±0.3
walker2d-med-exp 107.5 113.0 110.1 108.8 109.6 110.1 112.4 ±0.6
antmaze-umaze 54.6 64.3 78.6 74.0 87.5 87.2 95.8 ±1.5
antmaze-umaze-div 45.6 60.7 71.4 84.0 62.2 69.2 78.8 ±5.8
antmaze-med-play 0.0 0.3 10.6 61.2 71.2 73.5 62.0 ±6.0
antmaze-med-div 0.0 0.0 3.0 53.7 70.0 67.8 43.8 ±5.4
antmaze-large-play 0.0 0.0 0.2 15.8 39.6 41 24.8 ±4.7
antmaze-large-div 0.0 0.0 0.0 14.9 47.5 47.3 26.2 ±5.1
kitchen-complete 65.0 - - 43.8 62.5 72.5 90.8 ±3.4
kitchen-partial 38.0 - - 49.8 46.3 73.8 74.6 ±0.9
kitchen-mixed 51.5 - - 51.0 51.0 54.6 62.5 ±4.1
Table 1: Average normalized scores on Gym locomotion, AntMaze and Kitchen tasks. The highest scores are in bold, and the standard deviation across six seeds is displayed as $\pm$ for MXQL.
@@
Dataset MXQL
n=4 n=8 n=12
Gym 86.5 86.4 86.5
AntMaze 41.8 54.4 55.3
Kitchen 76.0 74.6 76.1
Table 3: Average scores of MXQL for various $n$ values in each domain of D4RL Offline RL.
@@ 5 Related Work
In Offline RL, some methods regularize through conservatism (Wu et al., 2019; Kumar et al., 2019; Fujimoto et al., 2019; Kumar et al., 2020; Fujimoto & Gu, 2021; Nair et al., 2021) and others that directly model the greedy policy (Peng et al., 2019; Brandfonbrener et al., 2021; Chen et al., 2021). XQL was able to learn without direct access to the current policy, similar to Kostrikov et al. (2022), while introducing conservatism. Similar to Online RL, methods for stabilization have been employed in Offline RL; notably, the stabilization in MXQL has shown performance improvements in certain Offline RL tasks. Recently, there have been methods that have led to performance improvements by using Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) as the policy (Wang et al., 2023; Hansen-Estruch et al., 2023), and applying this approach in MXQL could be a future direction.
@@ A.2 Offline RL
In Offline RL, the official implementation is also used, and the hyperparameters are the same as in (Garg et al., 2023). In Gym tasks, "-v2" was used, while in AntMaze and Kitchen tasks, "-v0" was employed. The batch size and the update frequency of the V-function, which were tuned in (Garg et al., 2023), are not tuned. In XQL, the $\beta$ and the size of clipping, and in MXQL, the $\beta$ and the order $n$ of expansion have been tuned, and a common value for each domain has been used. The range for tuning $\beta$ is the same as for (Garg et al., 2023), which is [0.6, 0.8, 1, 2, 5]. These values are shown in Table 5. The average scores for other $n$ in MXQL are shown in Table 6. The experiments are conducted using 6 random seeds.
@@
Dataset MXQL
n=4 n=8 n=12
halfcheetah-med 47.8 $\pm$ 0.1 47.8 $\pm$ 0.1 47.8 $\pm$ 0.1
hopper-med 90.0 $\pm$ 2.7 90.3 $\pm$ 3.0 89.7 $\pm$ 3.6
walker2d-med 85.9 $\pm$ 0.5 86.0 $\pm$ 0.4 86.1 $\pm$ 0.2
halfcheetah-med-rep 45.8 $\pm$ 0.1 45.7 $\pm$ 0.2 45.7 $\pm$ 0.3
hopper-med-rep 101.8 $\pm$ 0.6 101.7 $\pm$ 1.0 101.8 $\pm$ 0.8
walker2d-med-rep 89.0 $\pm$ 1.5 89.3 $\pm$ 1.8 88.8 $\pm$ 1.5
halfcheetah-med-exp 93.0 $\pm$ 0.3 93.0 $\pm$ 0.5 93.3 $\pm$ 0.7
hopper-med-exp 112.4 $\pm$ 1.8 111.8 $\pm$ 1.7 113.0 $\pm$ 0.3
walker2d-med-exp 112.7 $\pm$ 1.5 111.9 $\pm$ 0.5 112.4 $\pm$ 0.6
antmaze-umaze 95.3 $\pm$ 0.9 94.8 $\pm$ 1.8 95.8 $\pm$ 1.5
antmaze-umaze-div 69.8 $\pm$ 4.4 81.5 $\pm$ 3.0 78.8 $\pm$ 5.8
antmaze-med-play 35.8 $\pm$ 4.6 66.7 $\pm$ 6.2 62.0 $\pm$ 6.0
antmaze-med-div 25.8 $\pm$ 3.4 47.7 $\pm$ 7.8 43.8 $\pm$ 5.4
antmaze-large-play 10.3 $\pm$ 2.9 19.5 $\pm$ 8.1 24.8 $\pm$ 4.7
antmaze-large-div 13.8 $\pm$ 3.7 16.3 $\pm$ 3.2 26.2 $\pm$ 5.1
kitchen-complete 90.8 $\pm$ 3.4 87.5 $\pm$ 2.0 90.0 $\pm$ 3.8
kitchen-partial 74.6 $\pm$ 0.9 73.8 $\pm$ 2.2 72.9 $\pm$ 2.2
kitchen-mixed 62.5 $\pm$ 4.1 62.5 $\pm$ 2.0 65.4 $\pm$ 5.7
Table 6: Average scores of MXQL for various values of $n$ in Offline RL tasks from D4RL.

### v2

@@ Abstract
In offline reinforcement learning, in-sample learning methods have been widely used to prevent performance degradation caused by evaluating out-of-distribution actions from the dataset. Extreme Q-learning (XQL) employs a loss function based on the assumption that Bellman error follows a Gumbel distribution, enabling it to model the soft optimal value function in an in-sample manner. It has demonstrated strong performance in both offline and online reinforcement learning settings. However, issues remain, such as the instability caused by the exponential term in the loss function and the risk of the error distribution deviating from the Gumbel distribution. Therefore, we propose Maclaurin Expanded Extreme Q-learning to enhance stability. In this method, applying Maclaurin expansion to the loss function in XQL enhances stability against large errors. This approach involves adjusting the modeled value function between the value function under the behavior policy and the soft optimal value function, thus achieving a trade-off between stability and optimality depending on the order of expansion. It also enables adjustment of the error distribution assumption from a normal distribution to a Gumbel distribution. Our method significantly stabilizes learning in online RL tasks from DM Control, where XQL was previously unstable. Additionally, it improves performance in several offline RL tasks from D4RL.
@@ 1 Introduction
Thus, in this study, we propose a both simple and practical algorithm, Maclaurin Expanded Extreme Q-learning (MXQL), which not only stabilizes the Gumbel loss but also allows for the adjustment of the error distribution assumption from normal to Gumbel. The proposed method uses Expanded Gumbel loss, which is a Maclaurin expansion of the Gumbel loss. By reducing the order of expansion $n$, this loss, as shown to the left of Figure 1, mitigates excessively large or small gradients compared to the Gumbel loss.
As $n$ increases, the loss function converges to the Gumbel loss, while for $n=2$, it becomes the L2 loss. With the Gumbel loss, the estimated value function becomes a soft optimal value, and with the L2 loss, it becomes the value function under the behavior policy. These learning methods correspond to soft Q-learning (Haarnoja et al., 2017) and SARSA, respectively. By adjusting $n$ between $2$ and $\infty$, the method of estimating the value function can be adjusted between soft Q-learning and SARSA. In other words, a larger $n$ leads to an unstable but optimal value function estimation, while a smaller $n$ results in a stable but non-maximizing estimation, offering a trade-off between stability and optimality. This is analogous to adjusting the parameter $\tau$ in IQL’s expectile loss from $1$ to a smaller value for stabilization. IQL adjusts between SARSA and Q-learning using the parameter $\tau$, while MXQL adjusts between SARSA and soft Q-learning using the parameter $n$.
The assumed error distribution follows a normal distribution when $n=2$ because the loss is the L2 loss. When $n$ is large, it follows a Gumbel distribution. Therefore, by adjusting $n$ between $2$ and a larger value, the assumed error distribution can be adjusted between the normal and Gumbel distributions. In practice, the Bellman error is influenced by various factors such as the target network and double Q-learning, which suggests that, according to the central limit theorem, it may approximate a normal distribution. The assumed error distribution is illustrated to the right of Figure 1, showing that as $n$ increases, the distribution transitions from normal to Gumbel.
In the experiments, similar to Garg et al. (2023), we compared performance using DM Control tasks (Tassa et al., 2018; Tunyasuvunakool et al., 2020) for online RL and D4RL tasks (Fu et al., 2020) for offline RL. In online RL scenarios, where XQL was previously unstable, we observed improved stability. Furthermore, in offline RL, our method demonstrated superior performance compared to existing methods, including XQL, across several tasks.
@@ 2.2.2 Gumbel Regression
In XQL, this property is applied to Q-learning, directly estimating the soft-value function ($V^{*}(s)=\beta\log\sum_{a}\mu(a\mid s)\text{exp}(Q(s,a)/\beta)$) in maximum entropy RL. The loss function becomes:
@@
The Q-function is learned using the least squares method as follows:
$L(Q)=\mathbb{E}_{s,a,s^{\prime}\sim\mathcal{D}}[(r(s,a)+\gamma V(s^{\prime})-Q(s,a))^{2}]$ (5)
@@ 2.2.3 Role of $\beta$
The soft optimal value estimated in Eq. 4 is derived from the following objective function in Maximum Entropy RL.
$J(\pi)=\mathbb{E}_{s,a\sim\pi}\left[Q(s,a)\right]-\beta D_{\text{KL}}(\pi||\mu)\$ (6)
Here, $\beta D_{\text{KL}}(\pi||\mu)$ serves as a conservative term against deviations from the behavior policy, with $\beta$ modulating the level of conservatism. As can be seen from Eq. 4, this parameter, $\beta$, also impacts learning stability: small $\beta$ values can increase gradients, risking divergence, while large $\beta$ values may reduce gradient magnitudes, decelerating convergence. Therefore, while $\beta$ modifies conservatism, it does not ensure stability across all values.
@@ 3.1 Instability related to $\beta$
For scalar data represented as $x_{i}$, parameter $h$, and Gumbel noise $\epsilon_{i}$, we perform Gumbel Regression using Eq. 3 as $x_{i}=h+\epsilon_{i}$ where $\epsilon_{i}\sim-\mathcal{G}(0,\beta_{reg}),x_{i}\sim-\mathcal{G}(0,\beta_{data})$. This implies that when $\beta_{reg}$ and $\beta_{data}$ are equal, the assumptions concerning the error distribution in Gumbel Regression are fully met. We varied the $\beta$ of the data ($\beta_{data}$) and the $\beta$ used in Gumbel Regression ($\beta_{reg}$) respectively, estimated them, and analyzed how differences between $\beta_{data}$ and $\beta_{reg}$ affected the estimation results. The estimated $h$ was assessed based on its deviation from the true value ($\log\sum e^{x_{i}/\beta_{reg}}$). The results are presented in Figure 2.
@@ 3.2 Expanded Gumbel Loss
where $n$ is the order of expansion. As shown on the left side of Figure 1, the Expanded Gumbel loss exhibits a less steep gradient on the right side and a steeper gradient on the left side compared to the Gumbel loss. This modification resolves issues such as the collapse of learning and slow convergence observed in the preliminary experiments. Moreover, by adjusting the expansion order $n$, the degree of stabilization can be altered: smaller values of $n$ lead to greater deviation from the Gumbel loss but enhance stability. It should be noted that $n$ is selected from even numbers to ensure the loss remains positive. In MXQL, similar to XQL, the Expanded Gumbel loss is employed for estimating the value function, and the loss function is defined as follows:
@@
Learning the Q-function uses Eq. 5 in the same manner as in XQL.
As shown in Eq. 8, when $n=2$, the loss function becomes the L2 loss. The V-function represents the expected value of Q for actions taken under the behavior policy. In other words, the learning of the value function follows a SARSA-based approach, and the estimated value function is the value under the behavior policy rather than the optimal policy. When $n$ approaches $\infty$, Eq. 8 transforms into the Gumbel loss. The value function estimated by the Gumbel loss corresponds to the soft optimal value in maximum entropy RL. Thus, the learning of the value function follows a soft Q-learning-based approach. This indicates a trade-off between stability and optimality: when $n$ is large, the estimated value function is more optimal but less stable, whereas when $n$ is small, the value function is more stable but but does not fully maximize.
This trade-off is similar to the parameter $\tau$ in IQL (Kostrikov et al., 2022), which adjusts the expectile loss. In IQL, when $\tau=0.5$, the loss becomes the L2 loss, corresponding to $n=2$ in MXQL, while $\tau=1$ in IQL corresponds to $n=\infty$ in MXQL. IQL adjusts between SARSA and Q-learning using the parameter $\tau$, while MXQL adjusts between SARSA and soft Q-learning using the parameter $n$. Since IQL typically uses values of $\tau$ less than $1$, such as $0.7$ or $0.9$, this suggests the necessity of stabilizing the loss in XQL towards the L2 loss.
@@ 3.3 Perspective of Error Distribution
We name our method, which substitutes XQL’s Gumbel loss with Expanded Gumbel loss, Maclaurin Expanded Extreme Q-learning (MXQL). Our approach is based on XQL, simply changing the loss function for V-function. The implementation details are provided in the appendix. In MXQL, the expansion order $n$ is a hyperparameter. However, it should be noted that the clipping size is also a hyperparameter in XQL, and the number of hyperparameters has not changed compared to XQL.
@@ 4.2 Offline RL
To ensure a fair comparison with XQL, experiments were conducted on the same Gym locomotion, AntMaze, and Franka Kitchen tasks as those reported in Garg et al. (2023). The scores are displayed in Table 1 and are compared with those from Brandfonbrener et al. (2021); Fujimoto & Gu (2021); Kumar et al. (2020); Kostrikov et al. (2022); Garg et al. (2023). The proposed method, MXQL, demonstrated the best scores in several tasks compared to the scores of existing methods reported in Kostrikov et al. (2022); sikchi2024dual, and outperformed XQL in most tasks. In the AntMaze medium and large tasks, MXQL underperforms compared to IQL, suggesting that the soft optimal value estimated in both XQL and MXQL may not be well-suited for these particular tasks.
@@
halfcheetah-med 42.6 48.4 48.3 44.0 47.4 47.4 46.5 $\pm$ 0.3
hopper-med 52.9 59.6 59.3 58.5 66.3 68.5 68.3 $\pm$ 7.3
walker2d-med 75.3 81.8 83.7 72.5 78.3 81.4 71.9 $\pm$ 3.9
halfcheetah-med-rep 36.6 38.1 44.6 45.5 44.2 44.1 44.1 $\pm$ 0.3
hopper-med-rep 18.1 97.5 60.9 95.0 94.7 95.1 98.2 $\pm$ 3.2
walker2d-med-rep 26.0 49.5 81.8 77.2 73.9 58.0 57.9 $\pm$ 8.8
halfcheetah-med-exp 55.2 93.4 90.7 91.6 86.7 90.8 88.2 $\pm$ 4.4
hopper-med-exp 52.5 103.3 98.0 105.4 91.5 94.0 105.1 $\pm$ 10.1
walker2d-med-exp 107.5 113.0 110.1 108.8 109.6 110.1 109.9 $\pm$ 0.1
antmaze-umaze 54.6 64.3 78.6 74.0 87.5 47.7 88.3 $\pm$ 2.1
antmaze-umaze-div 45.6 60.7 71.4 84.0 62.2 51.7 53.2 $\pm$ 9.7
antmaze-med-play 0.0 0.3 10.6 61.2 71.2 31.2 50.8 $\pm$ 2.7
antmaze-med-div 0.0 0.0 3.0 53.7 70.0 0.0 52.2 $\pm$ 6.6
antmaze-large-play 0.0 0.0 0.2 15.8 39.6 10.7 18.7 $\pm$ 5.9
antmaze-large-div 0.0 0.0 0.0 14.9 47.5 31.3 14.3 $\pm$ 6.4
kitchen-complete 65.0 - - 43.8 62.5 56.7 64.2 $\pm$ 10.3
kitchen-partial 38.0 - - 49.8 46.3 48.6 47.1 $\pm$ 8.7
kitchen-mixed 51.5 - - 51.0 51.0 40.4 71.9 $\pm$ 3.6
Table 1: Average normalized scores on Gym locomotion, AntMaze and Kitchen tasks. Highlighted results are within one performance point of those achieved by the best-performing algorithm, and the standard deviation across six seeds is displayed as $\pm$ for MXQL.
@@ 5 Related Work
In offline RL, some methods regularize through conservatism (Wu et al., 2019; Kumar et al., 2019; Fujimoto et al., 2019; Kumar et al., 2020; Fujimoto & Gu, 2021; Nair et al., 2021) and others that directly model the greedy policy (Peng et al., 2019; Brandfonbrener et al., 2021; Chen et al., 2021). XQL was able to learn without direct access to the current policy, similar to Kostrikov et al. (2022); xu2023sqleql, while introducing conservatism. Similar to online RL, methods for stabilization have been employed in offline RL; notably, the stabilization in MXQL has shown performance improvements in certain offline RL tasks. Recently, there have been methods that have led to performance improvements by using Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) as the policy (Wang et al., 2023; Hansen-Estruch et al., 2023), and applying this approach in MXQL could be a future direction.
@@ A.2 Offline RL
In offline RL, the official implementation is also used, and the hyperparameters are the same as in Garg et al. (2023). In Gym tasks, "-v2" was used, while in AntMaze and Kitchen tasks, "-v0" was employed. The batch size and the update frequency of the V-function, which were tuned in (Garg et al., 2023), are not tuned. In XQL, the $\beta$ and the size of clipping, and in MXQL, the $\beta$ and the order $n$ of expansion have been tuned, and a common value for each domain has been used. The range for tuning $\beta$ is the same as for Garg et al. (2023), which is [0.6, 0.8, 1, 2, 5]. The $n$ was selected from $[4,8,12,16,20]$. These values are shown in Table 4. In Garg et al. (2023), the evaluation was based on the best score during the learning process rather than the final score, and therefore, the XQL scores in Table 1 are cited from sikchi2024dual. The experiments are conducted using 6 random seeds.